The XYZ Ltd. contains the following information:
Balance sheet as on March 31, 2017
Particulars Previous year (Rs. thousand) Current year (Rs thousand)
Cash 200 150
Sundry debtors 320 400
Temporary investments 200 320
Stock 1840 2160
prepaid expenses 28 12
Total current assets 2,588 3,052
Total assets 5,600 6,400
Current liabilities 640 800
5% Debentures 1,600 1,600
Equity share capital 2,000 2,000
Retained earnings 468 904
Statement of profits year ended March 31, 2017
Particulars (Rs thousand)
Sales 4000
Less: Cost of goods sold 2800
Less: Interest 160
Net profit for current year 1040
Less: Taxes 364
Earnings after taxes 676
Dividend declared on equity shares 220

From the above, appraise the financial position of the company from
1
Expert's answer
2019-10-29T10:08:25-0400

1. The ratio of working capital = 3052 /800 = 3.815 — pretty sound.

The working capital ratio is calculated by dividing current assets by current liabilities.

2.Earnings per share

Buying stocks, you participate in the future profits (or the risk of loss) of the company. Earnings per share (EPS) measures the net profit obtained in each shares the common shares of the company. Analysts company divided its net income by the weighted average number of ordinary shares that have come out during the year.

3.The coefficient value and wages

In short, this ratio reflects the investors' assessment of future income. You determine the price of shares of the company and divide it into EPS, to get the ratio of P / E.

4.debt and capital ratio

What to do if your prospective investment objective of taking too much? This may reduce the margin of safety for the fact that he has to raise the fixed charges, reduce profits for dividends for the likes of you, and even trigger a financial crisis.

The loan to equity ratio is calculated by adding the outstanding long and short-term liabilities, and dividing by the book value of equity.

5.Return on equity = (1040-220)/2000 = 0.41

Common shareholders want to know how profitable their capital in companies in which they invest. Return on equity is calculated by taking the company's net profit (after tax) by subtracting preferred dividends and dividing the result by the total dollar company.

This gives ROE 41%. The higher the ROE, the better the company makes a profit.



https://www.investopedia.com/articles/fundamental/04/063004.asp

https://www.investopedia.com/articles/04/031004.asp

https://www.investopedia.com/financial-edge/0910/6-basic-financial-ratios-and-what-they-tell-you.aspx

https://www.afponline.org/ideas-inspiration/topics/articles/Details/6-steps-to-an-effective-financial-statement-analysis